If producing 200 buttons and 200 safety pins daily is a 50 1/0 split of resources, where do we see the opportunity cost if you decide to produce 300 buttons and 100 safety pins ?
1
Expert's answer
2020-11-25T12:16:48-0500

If you increase the production of buttons from 200 to 300, you need to reduce the production of pins from 200 to 100. That is, increasing the production of buttons by 100 units requires reducing pins by 100 units.

And according to the law of increasing opportunity costs: in order to get more of one type of good in a given time period, society must sacrifice an ever-increasing amount of another good:

"opportunity cost=\frac{200-100}{300-200}=\frac{100}{100}=1"